# UStyle: Waterbody Style Transfer of Underwater Scenes by Depth-Guided Feature Synthesis

### v1

@@ V-C Quantitative Results
We conduct a comprehensive quantitative evaluation using five complementary metrics across various underwater STx scenarios from the UF7D test set. As presented in Table I(a), UStyle achieves 4×4$\times$4 × to 6×6$\times$6 × lower RMSE rates across all scenarios compared to SOTA artistic and photorealistic STx methods. It also consistently outperforms competing approaches in PSNR and SSIM, as shown in Table I(b). The closest competitor, AquaFuse, requires a carefully selected reference image and prior information such as incident angle and attenuation characteristics. In contrast, UStyle achieves improvements of $6.8\%$ (RMSE), $2.1\%$ (PSNR), and $3.3\%$ (SSIM) over AquaFuse while operating without any reference or prior knowledge — demonstrating its robustness and accuracy for waterbody STx.
@@ V-D Ablation Studies
VI Analysis and Discussion
VI-A Analysis of Results

### v2

@@ V-C Quantitative Results
We conduct a comprehensive quantitative evaluation using five complementary metrics across various underwater STx scenarios from the UF7D test set. As presented in Table I(a), UStyle achieves 4×4$\times$4 × to 6×6$\times$6 × lower RMSE rates across all scenarios compared to SOTA artistic and photorealistic STx methods. It also consistently outperforms competing approaches in PSNR and SSIM, as shown in Table I(b). The closest competitor, AquaFuse, requires a carefully selected reference image and prior information such as incident angle and attenuation characteristics. In contrast, UStyle achieves improvements of $15\%$ (RMSE), $5.29\%$ (PSNR), and $5.71\%$ (SSIM) over AquaFuse while operating without any reference or prior knowledge — demonstrating its robustness and accuracy for waterbody STx.
@@ V-D Ablation Studies
V-E Analysis of Results
@@
Figure 11: Evaluation of UStyle’s generalization capability under three unseen scenarios: (a) Style transfer on previously unseen color hues (yellow and grey) from UIEB [123], (b) Consistent interpolation across progressively intensified waterbody styles from the TURBID [124] dataset, (c) Illumination of low-light underwater scenes from CaveSeg [125] and OceanDark [126] datasets using a well-lit image as style.
VI Generalization Performance
VI-A Unseen Style Transfer
The generalization ability of UStyle is validated through three distinct scenarios as shown in Figure 11, using the model trained on the UF7D dataset without any additional fine-tuning. First, in Figure 11 (a) , We used UIEB [123], which feature unseen waterbody hues such as grey and yellow, distinct from the predominantly blue and green hues seen during training. UStyle demonstrated its ability to transfer these unseen hues onto the content images, highlighting its strong generalization capability and robust style representation learned from the UF7D dataset. Next, we examine gradual style intensification using the TURBID [124] dataset in Figure 11 (b). This dataset provides images that are originally clear, along with corresponding versions exhibiting progressively increasing levels of turbidity (light to dark blue tone). Using these as style references, we observe that UStyle performs smooth and consistent style transfer across all levels of distortion, indicating its ability to interpolate within a continuous style manifold.
Figure 11 emphasized the practical significance of UStyle by evaluating style transfer in low-light underwater conditions using images from the CaveSeg [125] and OceanDark [126] datasets. We applied a well-lit, high-quality style image to dimly illuminated underwater scenes. UStyle effectively transferred the brighter illumination conditions from the reference style image, enhancing visibility and clarifying structural details. This capability demonstrates a meaningful practical application, showcasing how UStyle can be effectively utilized for other underwater imaging tasks.